# SEEA-R1: Tree-Structured Reinforcement Fine-Tuning for Self-Evolving Embodied Agents

### v1

@@ Abstract
00footnotetext: * Equal contribution.00footnotetext: $\dagger$ Project leader: jason.ju@x-humanoid.com00footnotetext: $\ddagger$ Corresponding authors: shanghang@pku.edu.cn, jian.tang@x-humanoid.com
[…]
Figure 1: SEEA-R1 self-evolves by planning with perception and tree-based search, refining actions via reward signals to solve tasks from high-level instructions in embodied environments.
@@ 1 Introduction
In this context, reinforcement fine-tuning (RFT) guo2025deepseek ; shao2024deepseekmath has emerged as a promising paradigm: By providing reward-based feedback, RFT enables agents to learn from trial-and-error trajectories and gradually refine their decision-making policies. However, bridging RFT to embodied agents introduces unique challenges that are absent in symbolic domains such as mathematics or code generation. In this work, we identify two major challenges for reinforce fine-tuning agents in embodied environments: 1) Lack of intermediate feedback in multi-step reasoning tasks: In embodied tasks with long horizons and delayed sparse rewards, it is difficult to assign credit to intermediate decisions, making it hard for RFT to effectively guide policy learning. 2) Poor generalization of handcrafted reward functions: Existing RFT pipelines typically rely on simulator-specific or task-specific reward signals that do not generalize to novel environments, making it hard for agents to self-improve across diverse tasks.
To address these challenges, we introduce SEEA-R1 (Self-Evolving Embodied Agents), the first framework to adapt RFT for training embodied agents capable of long-horizon reasoning and autonomous self-evolution. SEEA-R1 integrates two key components. 1) Tree-GRPO (Tree-based Group Relative Policy Optimization): we extend Group Relative Policy Optimization with Monte Carlo Tree Search (MCTS), enabling agents to explore alternative trajectories and convert sparse outcome rewards into structured, step-wise process signals. This improves credit assignment and facilitates reasoning over extended action sequences. 2) MGRM (Multi-modal Generative Reward Model): To eliminate the reliance on handcrafted or environment-specific reward functions, we introduce MGRM—a reward model trained from multi-modal, multi-turn trajectories. MGRM provides task-agnostic reward estimation, supporting generalization and self-improvement across diverse embodied scenarios.
We evaluate SEEA-R1 on the ALFWorld benchmark, which rigorously tests an agent’s planning and reasoning capabilities by requiring it to map abstract goals to visually grounded action sequences. Our proposed method achieves state-of-the-art success rates of 85.07% and 36.19% on textual and multi-modal tasks, respectively, outperforming previous models including Qwen2.5-VL and GPT-4o. To further evaluate self-evolution in a realistic yet challenging setting, we replace ground-truth rewards with MGRM-based self-supervised signals; SEEA-R1 still achieves 80.30% and 23.88% on textual and multi-modal tasks, respectively, surpassing all open-source baselines. These results highlight the effectiveness and scalability of SEEA-R1 as a self-evolving embodied agent.
@@ 2.1 Self-Evolution Learning
Recent research has explored how large language and vision-language models can self-evolve through mechanisms such as self-play, reinforcement learning, and trajectory refinementxu2024survey . Welcome to the Era of Experience silver2025welcome highlights the importance of learning from experience and real-world interactions, emphasizing two key components: reward mechanism for environmental interaction and long-horizon perception-action task loop. SPC chen2025spcevolvingselfplaycritic utilizes an adversarial self-play environment to enhance reasoning, while rStar-Math guan2025rstar and Agent-Q putta2024agentqadvancedreasoning demonstrates that small models can improve using MCTS with reward modeling. Dreamsmooth lee2023dreamsmooth and R2I samsami2024mastering incorporate planning and prediction modeling into reward design to address long-horizon tasks. RAGEN wang2025ragen introduces a modular system to enhance LLM reasoning capabilities in multi-turn interactive environments. Building on these foundations, our approach integrates MCTS into the self-evolution loop, enabling action rollout, trajectory filtering, and policy refinement. We scale it for the first time to vision-language decision-making in long-horizon, physically grounded tasks.
@@ 2.2 Planning for Embodied Agent
LLM-Planner song2023llm , LLM+P liu2023llm+ , and RegionFocus luo2025visualtesttimescalinggui have demonstrated that large language and vision-language models can significantly enhance task performance. Planning is also critical for embodied agents to make sequential decisions and achieve long-term goals. Training-free approaches such as PIVOT nasiriany2024pivotiterativevisualprompting and ECoT zawalski2024robotic leverage the spatial reasoning capabilities of VLMs to plan effectively, leading to improved task success rates in embodied agents. In contrast, training-based methods like VIPER salim2025viper and MPO xiong2025mpo combine vision-language perception with LLM-based reasoning or meta-planning, but often depend on predefined modules or expert-designed strategies. Other works such as ARMAP chen2025scaling and LS-Imagine li2024open show that incorporating the planning module into the reward assignment process during training can significantly improve long-horizon task performance in web search and Minecraft environments. In the embodied domain, EmbodiedEval cheng2025embodiedeval highlights persistent planning failures in current embodied systems, indicating a gap between perception, reasoning, and decision execution. To address this, our approach integrates MCTS-based planning directly into training in embodied scenarios, which, to our knowledge, has not been previously explored in this context.
@@ 2.3 Reinforcement Fine-Tuning
DeepSeek-R1 guo2025deepseek  muennighoff2025s1simpletesttimescaling shows that RL with formatting and result-only rewards can steer LLMs toward humanlike, complex chain-of-thought reasoning, boosting performance on challenging tasks. Subsequently, Search-R1 jin2025search , TORL li2025torl explored extending the LLM-R1 approach to tool use. R1 approach has also achieved success in the visual domain. Vision-R1 huang2025vision , Visual-RFT liu2025visual , and Video-R1 zhan2025vision improve visual grounding through reward-based reinforcement fine-tuning, while R1-OneVision yang2025r1 and Perception-R1 yu2025perception formalize visual input into language for structured reasoning. In the embodied AI domain, robotxR1 boyle2025robotxr1 transforms visual information into textual descriptions to train LLMs, enhancing their robotic scenario decision-making capabilities. Embodied-R zhao2025embodied further extends to spatial reasoning. But they are not all trained as multimodal models and still rely on task-specific priors. To address these limitations, we developed the first general embodied R1 framework SEEA-R1, which integrates a vision-grounded reward model with MCTS-based planning to enable self-supervised policy optimization without handcrafted rewards.
Figure 2: SEEA-R1 framework. The framework drives continuous improvement through an iterative loop of two core cycles as follows: 1.Data Evolution: The Policy Model interacts with the environment via MCTS from an initial state to generate the experience dataset, containing trajectories with derived Q-values, ground truth rewards from the environment, and rewards from the current Reward Model. 2.Model Evolution: The collected data is used to update both models: (a) The Policy Model to predict actions and (b) The Reward Model to predict categorical outcomes. Refined models from Model Evolution then drive the next Data Evolution iteration, enabling continuous self-evolution.
@@ 3.2.2 Model Evolution: Co-refining Policy Model and Reward Model
[…]
@@ 3.2.3 Iterative Self-Evolution Loop
The SEEA-R1 framework iteratively alternates between Data Evolution and Model Evolution. In each cycle, the updated Embodied Agent and MGRM from the Model Evolution phase are deployed in the subsequent Data Evolution phase. This iterative evolution process, detailed in Appendix, allows the SEEA-R1 agent to progressively enhance its reasoning, reward understanding, and task execution capabilities over time.
@@ 4.1.1 Environment, Dataset and Evaluation
ALFWorld. We train and evaluate our approach on ALFWorld shridhar2020alfworld , an interactive simulation environment and a widely used benchmark for evaluating embodied agents on complex, long-horizon tasks in simulated household environments. We leverage this interactive benchmark to collect data for model training, using the rewards provided by ALFWorld as our ground truth (GT) rewards. ALFWorld provides two types of observations: textual descriptions and raw visual inputs (images), enabling evaluation in both language-only and vision-language settings. Further details on the dataset and expert trajectory collection can be found in Appendix.
EmbodiedEval. To evaluate the generalization ability of our embodied agents beyond the training environment, we introduce EmbodiedEval cheng2025embodiedeval as an out-of-distribution benchmark. EmbodiedEval tests MLLMs as embodied agents across diverse tasks—including Attribute Question Answering (AttrQA), Spatial Question Answering (SpatialQA), Navigation, Object Interaction, and Social Interaction—within 125 realistic 3D scenes. It provides a comprehensive assessment of agent capabilities in previously unseen scenarios. This setup enables us to measure generalization under significant domain shifts compared to the ALFWorld environment. We analyze the impact of different control strategies on performance, using key metrics such as overall accuracy, which reflects the percentage of fully completed tasks.
@@
To ensure experimental reproducibility, we set the decoding temperature to zero. When generating task completion trajectories, we include a one-shot in-context example for each task. Detailed prompts are provided in Appendix. Our primary evaluation metric is the Average Success Rate, which calculates the mean Success Rate across all test set task instances.
@@ 4.1.2 Implementation Details
MCTS
SEEA-R1 is employed to sample experience trajectories with process rewards for training. Each MCTS performs 30 search iterations, with each simulation proceeding up to a maximum depth of 30 steps. During the expansion phase, up to five child nodes—corresponding to candidate actions—are expanded from each selected leaf node.
Training Settings
We use the Qwen2.5-VL-7B-Instruct qwen2.5-VL as the base model to build our embodied agents. Since experience learning in our framework consists of two stages—trajectory sampling and model training—we define each iteration as one complete cycle of these two steps. To ensure fair comparison across different training paradigms (SFT, DPO and GRPO), we use consistent sample data sizes and training hyperparameters. For each iteration, the model is trained on newly collected data for one epoch as steps is 4. The total number of iterations is not fixed and proceeds until convergence. Specifically, for Tree-GRPO, model updates are performed after collecting 512 valid samples (i.e., with non-zero advantage values). For SFT, 512 trajectories with positive advantages are used. For DPO, 512 positive-negative trajectories pairs are sampled from the same parent node in the MCST. When training the multimodal generative reward model (MGRM) via GRPO, the group size is set to 24. All training adopt the same hyperparameters: a cosine annealing learning rate schedule (initial LR: 1e-6, warmup ratio: 0.05), batch size of 128, and KL divergence coefficient $\beta$ of 0.01. Experiments are conducted on 8 NVIDIA A100 80GB GPUs using the ms-swift framework zhao2024swiftascalablelightweightinfrastructure .
4.2 Experimental Results
4.2.1 Overall Performance Comparison
@@
SEEA-R1 7B Tree-GRPO 49 40 43 41 24 16 36
Our proposed SEEA-R1 method, utilizing Qwen2.5-VL-7B-Instruct and Qwen2.5-7B-Instruct as respective base models, was evaluated against strong baselines on ALFWorld test set. As reported in Tables 1 and 2, SEEA-R1 was tested in two settings: text+visual (MLLM), with image observations, and text-only (LLM), with textual descriptions. In the more challenging MLLM setting requiring direct visual perception, SEEA-R1 achieved a 36.19% success rate, substantially outperforming GPT-4o and other leading open-source models, and demonstrated superior efficiency (23.37 vs. 25.10 average steps for GPT-4o). Furthermore, in the text-only (LLM) setting, SEEA-R1 attained an 84.3% success rate, with average completion steps (14.75) on par with GPT-4o.
Table 2: Performance comparison on the ALFWorld dataset. ITO stands for Inference-Time Optimization. MPO (Meta Plan Optimization) xiong2025mpo provides the agent with the meta plan as context in inference time to improve performance.
Model w/o Training ITO ALFWorld Average
Seen Unseen
GPT-4o* achiam2023gpt - 78.6 83.6 81.1
Qwen2.5-7B-Instruct* Yang2024Qwen25TR 71.4 75.4 73.4
Llama-3.1-8B-Instruct* grattafiori2024llama - 22.9 28.4 25.7
Llama-3.1-70B-Instruct* grattafiori2024llama - 78.6 73.9 76.3
GPT-4o* MPO 89.3 93.3 91.3
Llama-3.1-8B-Instruct* MPO 50.0 52.2 51.1
Llama-3.1-8B-Instruct + SFT* zeng2023agenttuning - 79.3 71.6 75.5
Llama-3.1-8B-Instruct + ETO* song2024trial - 77.1 76.4 76.8
Llama-3.1-8B-Instruct + KnowAgent* zhu2024knowagent - 80.0 74.9 77.5
Llama-3.1-8B-Instruct + WKM* qiao2024agent - 77.1 78.2 77.7
Llama-3.1-8B-Instruct + SFT* MPO 80.7 81.3 81.0
Llama-3.1-8B-Instruct + ETO* MPO 85.0 79.1 82.1
SEEA-R1 - 85.3 85.1 85.2

### v2

@@ Abstract
Self-evolution, the ability of agents to autonomously improve their reasoning and behavior, is essential for the embodied domain with long-horizon, real-world tasks. Despite current advancements in reinforcement fine-tuning (RFT) showing strong performance in enhancing reasoning in LLMs, its potential to enable self-evolving embodied intelligence with multi-modal interactions remains largely unexplored. Specifically, reinforcement fine-tuning faces two fundamental obstacles in embodied settings: (i) the lack of accessible intermediate rewards in multi-step reasoning tasks limits effective learning signals, and (ii) reliance on hand-crafted reward functions restricts generalization to novel tasks and environments. To address these challenges, we present SSSSSSSSSSSSSSSSSelf-EEEEEEEEEEEEEEEEEvolving EEEEEEEEEEEEEEEEEmbodied AAAAAAAAAAAAAAAAAgents-R1, SEEA-R1, the first RFT framework designed for enabling the self-evolving capabilities of embodied agents. Specifically, to convert sparse delayed rewards into denser intermediate signals that improve multi-step reasoning, we propose Tree-based Group Relative Policy Optimization (Tree-GRPO) integrates Monte Carlo Tree Search into GRPO. To generalize reward estimation across tasks and scenes, supporting unsupervised adaptation and reward-driven self-evolution, we further introduce the Multi-modal Generative Reward Model (MGRM). To holistically evaluate the effectiveness of SEEA-R1, we evaluate on the ALFWorld benchmark, surpassing state-of-the-art methods with scores of 85.07% (textual) and 46.27% (multi-modal), outperforming prior models including GPT-4o. SEEA-R1 also achieves scores of 80.3% (textual) and 44.03% (multi-modal) without ground truth reward, surpassing all open-source baselines and highlighting its scalability as a self-evolving embodied agent. Additional experiments and qualitative analysis further support the potential of SEEA-R1 for future research in scalable embodied intelligence. Project page is at https://seea-r1.github.io/.
Figure 1: SEEA-R1 self-evolves by reasoning over its environment with perception-grounded planning. The agent explores task solutions using tree-based search guided by a reward model, iteratively refining actions to achieve complex goals. Given a high-level instruction, it explores, plans, and executes actions in an embodied environment.
@@ 1 Introduction
In this context, reinforcement fine-tuning (RFT) guo2025deepseek ; shao2024deepseekmath has emerged as a promising paradigm: By providing reward-based feedback, RFT enables agents to learn from trial-and-error trajectories and gradually refine their decision-making policies. However, bridging RFT to embodied agents introduces unique challenges that are absent in symbolic domains such as mathematics or code generation. In this work, we identify two major challenges for reinforcement fine-tuning agents in embodied environments: 1) Lack of intermediate feedback in multi-step reasoning tasks: in embodied tasks with long horizons and delayed sparse rewards, it is difficult to assign credit to intermediate decisions, making it hard for RFT to guide policy learning effectively. 2) Poor generalization of handcrafted reward functions: existing RFT pipelines typically rely on simulator-specific or task-specific reward signals that do not generalize to novel environments, harming agents’ abilities to self-improve across diverse tasks.
To address these challenges, we introduce SEEA-R1 (Self-Evolving Embodied Agents), the first framework to adopt RFT for training embodied agents capable of long-horizon reasoning and autonomous self-evolution. SEEA-R1 integrates two key components. 1) Tree-GRPO (Tree-based Group Relative Policy Optimization): we extend Group Relative Policy Optimization with Monte Carlo Tree Search (MCTS), enabling agents to explore alternative trajectories and convert sparse outcome rewards into structured, step-wise process signals. This improves credit assignment and facilitates reasoning over extended action sequences. 2) MGRM (Multi-modal Generative Reward Model): To eliminate the reliance on handcrafted or environment-specific reward functions, we introduce MGRM—a reward model trained from multi-modal, multi-turn trajectories. MGRM provides task-agnostic reward estimation, supporting generalization and self-improvement across diverse embodied scenarios. Moreover, SEEA-R1 utilizes a novel joint training paradigm: better MLLM reasoning enhances MGRM accuracy, while refined MGRM rewards further advance MLLM policy learning, achieving human-free, self-evolving embodied intelligence.
We evaluate SEEA-R1 on the ALFWorld benchmark, which rigorously tests an agent’s planning and reasoning capabilities by requiring it to map abstract goals to visually grounded action sequences. Our proposed method achieves state-of-the-art success rates of 85.07% and 46.27% on textual and multi-modal tasks, respectively, outperforming previous models including Qwen2.5-VL and GPT-4o. To further evaluate self-evolution in a realistic yet challenging setting, we replace ground-truth rewards with MGRM-based self-supervised signals; SEEA-R1 still achieves 80.30% and 44.03% on textual and multi-modal tasks, respectively, surpassing all open-source baselines. These results highlight the effectiveness and scalability of SEEA-R1 as a self-evolving embodied agent.
@@ 2.1 Self-Evolution Learning
Recent research has explored how large language and vision-language models can self-evolve through mechanisms such as self-play, reinforcement learning, and trajectory refinementxu2024survey . Welcome to the Era of Experience silver2025welcome highlights the importance of learning from experience and real-world interactions, emphasizing two key components: reward mechanism for environmental interaction and long-horizon perception-action task loop. SPC chen2025spcevolvingselfplaycritic utilizes an adversarial self-play environment to enhance reasoning, while rStar-Math guan2025rstar and Agent-Q putta2024agentqadvancedreasoning demonstrates that small models can improve using MCTS with reward modeling. Dreamsmooth lee2023dreamsmooth and R2I samsami2024mastering incorporate planning and prediction modeling into reward design to address long-horizon tasks. RAGEN wang2025ragen introduces a modular system to enhance LLM reasoning capabilities in multi-turn interactive environments. Building on these foundations, our approach integrates MCTS into the self-evolution loop, enabling action rollout, trajectory filtering, and policy-reward co-refinement.
@@ 2.2 Planning for Embodied Agent
LLM-Planner song2023llm , LLM+P liu2023llm+ , and RegionFocus luo2025visualtesttimescalinggui have demonstrated that large language and vision-language models can significantly enhance task performance. Planning is also critical for embodied agents to make sequential decisions and achieve long-term goals. Training-free approaches such as PIVOT nasiriany2024pivotiterativevisualprompting and ECoT zawalski2024robotic leverage the spatial reasoning capabilities of VLMs to plan effectively, leading to improved task success rates in embodied agents. In contrast, training-based methods like VIPER salim2025viper and MPO xiong2025mpo combine vision-language perception with LLM-based reasoning or meta-planning, but often depend on predefined modules or expert-designed strategies. ARMAP chen2025scaling and LS-Imagine li2024open show that incorporating the planning module into the reward assignment process during training can significantly improve long-horizon task performance in web search and Minecraft environments. Wang et al. Wang2025WorldMM combines tree search with preference optimization to improve state prediction and action selection, boosting embodied planning performance.
@@ 2.3 Reinforcement Fine-Tuning
DeepSeek-R1 guo2025deepseek  muennighoff2025s1simpletesttimescaling shows that RL with formatting and result-only rewards can steer LLMs toward humanlike, complex chain-of-thought reasoning, boosting performance on challenging tasks. Subsequently, Search-R1 jin2025search , TORL li2025torl explored extending the LLM-R1 approach to tool use, and AlphaLLM-CPL wang2024selfimprovementllmsmctsleveraging utilizes MCTS for self-improving preference learning. The R1 approaches have also achieved success in the visual domain. Vision-R1 huang2025vision , Visual-RFT liu2025visual , and Video-R1 zhan2025vision improve visual grounding through reward-based reinforcement fine-tuning, while R1-OneVision yang2025r1 and Perception-R1 yu2025perception formalize visual input into language for structured reasoning. In the embodied AI domain, robotxR1 boyle2025robotxr1 transforms visual information into textual descriptions to train LLMs, enhancing their robotic scenario decision-making capabilities. Embodied-R zhao2025embodied further extends to spatial reasoning. However, these methods are not trained as multimodal models and still rely on task-specific priors. To address these limitations, we developed the first general embodied RL framework SEEA-R1, which integrates a vision-grounded reward model with MCTS-based planning to enable self-supervised policy optimization without handcrafted rewards.
Figure 2: SEEA-R1 framework. The framework drives continuous improvement through an iterative loop of two core cycles as follows: 1.Data Evolution: The Policy Model interacts with the environment via MCTS from an initial state to generate the experience dataset, containing trajectories with derived Q-values, ground truth rewards from the environment, and rewards from the current Reward Model. 2.Model Evolution: The collected data is used to update both models: (a) Policy Model to predict actions and (b) Reward Model to predict categorical outcomes. Refined models from Model Evolution then drive the next Data Evolution iteration, enabling continuous self-evolution.
@@ 3.2.2 Model Evolution: Co-refining Policy Model and Reward Model
To improve the generalizability of the reward function, we adopt a learned reward model as an alternative to handcrafted reward. We design the reward model with several key properties: multi-modal for perception, multi-round for sequential interactions, and high interpretability. Specifically, our reward model is built upon a Multi-modal Large Language Model (MLLM). It takes the historical context $s_{t}$ as input and predicts one of three categorical outcomes: success, continue, or failure, whose prompt is in Appendix N.4. This allows us to leverage the DeepSeek-R1-Zero guo2025deepseek training paradigm to employ the GRPO shao2024deepseekmath for reinforcement learning , supporting two training paradigms tailored to scenarios with or without ground truth (GT) rewards.
With GT (Supervised Paradigm): MGRM first undergoes supervised fine-tuning (SFT) on GT-aligned trajectories (e.g., ALFWorld task completion/step validity signals), minimizing cross-entropy loss between predicted outcomes and GT labels to achieve sufficient performance and eliminate initial biases. During co-training with the policy, MGRM continues supervised training by calibrating against GT every 500 episodes—preventing drift and keeping judgments aligned with the simulator’s objective truth.
Without GT (Self-supervised Paradigm): In GT-free scenarios (e.g., real-world environments), MGRM relies on Test test Reinforcement Learning (TTRL) Zuo2025TTRLTR and GRPO: the policy generates K (set 10 in environments) diverse trajectories per initial state $s_{0}$ (via MCTS exploration), and MGRM’s majority-voted predictions across these trajectories form pseudo-GT for $s_{0}$. MGRM is then trained via GRPO (reward = +1 for matching pseudo-GT, 0 otherwise).
@@ 3.2.3 Iterative Self-Evolution Loop
The SEEA-R1 framework iteratively alternates between Data Evolution and Model Evolution. In each cycle, the updated Embodied Agent and MGRM from the Model Evolution phase are deployed in the subsequent Data Evolution phase. This iterative evolution process allows the SEEA-R1 agent to progressively enhance its reasoning, reward understanding, and task execution capabilities over time.
@@ 4.1.1 Environment, Dataset and Evaluation
ALFWorld. We train and evaluate our approach on ALFWorld shridhar2020alfworld , an interactive simulation environment and a widely used benchmark for evaluating embodied agents on complex, long-horizon tasks in simulated household environments. We leverage this interactive benchmark to collect data for model training, using the rewards provided by ALFWorld as our ground truth (GT) rewards. ALFWorld provides two types of observations: textual descriptions and raw visual inputs (images), enabling evaluation in both language-only and vision-language settings. Further details on the dataset and expert trajectory collection can be found in Appendix  F and  H.
EmbodiedEval.
To evaluate the generalization ability of SEEA-R1 beyond the training environment, we introduce EmbodiedEval cheng2025embodiedeval as an out-of-distribution benchmark. EmbodiedEval tests MLLMs as embodied agents across diverse tasks, including Attribute Question Answering (AttrQA), Spatial Question Answering (SpatialQA), Navigation, Object Interaction (Obj), and Social Interaction (Social), within 125 realistic 3D scenes. It provides a comprehensive assessment of agent capabilities in previously unseen scenarios. This setup enables us to measure generalization under significant domain shifts compared to the ALFWorld environment. We analyze the impact of different control strategies on performance, using key metrics such as overall accuracy, which reflects the percentage of fully completed tasks.
Real-World.
To evaluate SEEA-R1’s real-world performance and analyze the sim2real gap impact on embodied planning, we conducted physical experiments using a dual-arm ARX LIFT 2 with Mobile ALOHA-style teleoperation Fu2024MobileAL . Following RoboVQA Sermanet2023RoboVQAML ’s protocol, agents received long-horizon instructions and visual observations (front-view camera) to generate real-time actions for human-teleoperated and single-trajectory execution. We tested 3 task types (Pick & Place, Pick Two & Place, Clean) across 6 environments: apartment, reception room, tea room, family living room, children’s room, and kitchen. Each environment contained 12 test cases.
@@
To ensure experimental reproducibility, we set the decoding temperature to zero. When generating task completion trajectories, we include a one-shot in-context example for each task. Detailed prompts are provided in Appendix N. Our primary evaluation metric is the Average Success Rate, which calculates the mean Success Rate across all test set task instances.
@@ 4.1.2 Implementation Details
SEEA-R1 (w/o GT reward) 7B Tree-GRPO 44 60 73 58 26 26 44
SEEA-R1 (w/ GT reward) 7B Tree-GRPO 43 42 60 41 29 40 46
Monte Carlo Tree Search (MCTS)
We uses two-level pruning to reduce complexity of MCTS: Probabilistic Expansion (50% chance to expand all $K$ actions per node, reducing redundancy) and Strict Path Budget (hard limit L=5L{=}5 on full $K$-expansions per path, constraining search space). Each MCTS runs 30 iterations with max simulation depth 30.
Training Setting.
We use the Qwen2.5-VL-7B-Instruct qwen2.5-VL as the base model to build our embodied agents. Since experience learning in our framework consists of two stages—trajectory sampling and model training—we define each iteration as one complete cycle of these two steps. To ensure fair comparison across different training paradigms (SFT, DPO and GRPO), we use consistent sample data sizes and training hyperparameters. For each iteration, the model is trained on newly collected data for one epoch as steps is 4. The total number of iterations is not fixed and proceeds until convergence. Specifically, for Tree-GRPO, model updates are performed after collecting 512 valid samples (i.e., with non-zero advantage values). For SFT, 512 trajectories with positive advantages are used. For DPO, 512 positive-negative trajectories pairs are sampled from the same parent node in the MCST. When training the multimodal generative reward model (MGRM) via GRPO, the group size is set to 10, which equals to the vote num using TTRL). All training adopt the same hyperparameters: a cosine annealing learning rate schedule (initial LR: 1e-6, warmup ratio: 0.05), batch size of 128, and KL divergence coefficient $\beta$ of 0.0. Experiments are conducted on 8 NVIDIA A100 80GB GPUs using the ms-swift framework zhao2024swiftascalablelightweightinfrastructure .
4.2 Experimental Results
4.2.1 Overall Performance Comparison
Our proposed SEEA-R1 method, utilizing Qwen2.5-VL-7B-Instruct and Qwen2.5-7B-Instruct as respective base models, was evaluated against strong baselines on ALFWorld test set. As reported in Tables 1 and 7, SEEA-R1 was tested in two settings: text+visual (MLLM), with image observations, and text-only (LLM), with textual descriptions. In the more challenging MLLM setting requiring direct visual perception, SEEA-R1 achieved a 46.27% success rate, substantially outperforming GPT-4o and other leading open-source models, and demonstrated superior efficiency (23.37 vs. 25.10 average steps for GPT-4o). Furthermore, in the text-only (LLM) setting, SEEA-R1 attained an 84.3% success rate, with average completion steps (14.75) on par with GPT-4o.
4.2.2 Real World Results
Figure 4: Visualization of SEEA-R1 executing the "put clothes into the washing machine" task in real-world settings, which demonstrates reflection-correction capability.
We evaluated SEEA-R1 via single-trajectory execution against the Qwen2.5-VL-7B-Instruct baseline in physical real-world scenarios. As shown in Figure  4, SEEA-R1 exhibits striking reflection and correction abilities: when an action fails (e.g., attempting to place clothes into a closed washing machine), it actively identifies the error, adjusts its behavior, and re-executes successfully—an ability fostered by our MCTS-based self-evolution training framework. Quantitatively, Table 2 reveals that SEEA-R1 achieves a 34.72% absolute improvement in overall success rate across all tasks and scenarios, with particularly notable gains in complex environments like kitchens (9/12 vs. 1/12) and children’s rooms (10/12 vs. 4/12). This performance leap stems from the robust feature learning enabled by MCTS during training: even without simulation-specific advantages (e.g., rewinds), the model generalizes its ability to anticipate long-term consequences to real-world dynamics. These results confirm that our approach not only transfers effectively to physical robot settings but also empowers models with enhanced adaptive decision-making through strengthened reflection-correction mechanisms.
Table 2: Performance Comparison on Real-World Tasks and Scenarios
Model Living Room Guest Room Apartment Children’s Room Kitchen Tea Room Overall
Qwen2.5-VL-7B-Instruct 58.3 50.0 25.0 33.3 8.3 25.0 33.3
SEEA-R1-7B (w/ GT reward) 83.3 58.3 58.3 83.3 75.0 50.0 68.1